If f(x)=5x+2 and g(x)=-2x find the following.
1. f(0)+2f(2)
2. f(g(-2))
Thank you for your time.
- Jennifer

Answer by stanbon(48558)

1. f(0) = 5(0)+2 = 2
f(2) = 5(2)+2 = 12
Therefore f(0)+ 2f(2) = 2 +2(12) = 26
2. f(g(-2}) = ?
g(-2) =-2(-2)=4
So, f(g(-2)) = f(4)=5(4)+2 = 22
Cheers,
Stan H.